# INSIGHT: A Survey of In-Network Systems for Intelligent, High-Efficiency AI and Topology Optimization

### v1

@@ 6.2 Enhanced Network Security through In-Network Intrusion Detection Systems
In-network computation is essential for improving network security by creating more efficient Intrusion Detection Systems (IDS) [43]. Without using the control plane, line-rate inference can be used to detect and mitigate security threats by integrating AI-based security solutions straight into programmable network devices [3]. This enables real-time network traffic analysis at wire speed, which helps identify anomalies and subtle attack patterns that conventional perimeter-based security solutions sometimes . Real-time intrusion detection benefits from the latency reduction. The reduction can be obtained by performing feature extraction at the data plane using languages such as P4 [3]. Hence, in-network IDS can effectively support real-time, wire-speed anomaly detection by performing feature extraction and AI-based inference directly at the data plane, thereby identifying subtle attack patterns with reduced latency without compromising network throughput.
@@ 6.3 Optimized Traffic Management and Quality of Service
AI-optimized in-network computation is essential for dynamic and intelligent traffic management and Quality of Service (QoS) mechanisms [43]. AI models can be embedded into network devices to make necessary decisions about resource allocation, traffic routing, and prioritization. These decisions are based on application requirements and real-time network conditions [34]. In addition, AI tools can be used to optimize forwarding rule storage, enhance content caching, and anticipate user needs for better load balancing [34]. To increase throughput and guarantee the required bandwidth, in-network ML can help with dynamic network path selection and service type prioritization. The ability to perform early-stage traffic classification in the data plane allows for the embedding of Service Level Objectives (SLOs) into packet headers, enabling stateless QoS provisioning [43].

### v2

@@ 6.2 Enhanced Network Security through In-Network Intrusion Detection Systems
In-network computation is essential for improving network security by creating more efficient Intrusion Detection Systems (IDS) [38]. Without using the control plane, line-rate inference can be used to detect and mitigate security threats by integrating AI-based security solutions straight into programmable network devices. This enables real-time network traffic analysis at wire speed, which helps identify anomalies and subtle attack patterns that conventional perimeter-based security solutions sometimes. Real-time intrusion detection benefits from the latency reduction. The reduction can be obtained by performing feature extraction at the data plane using languages such as P4 [32]. Hence, in-network IDS can effectively support real-time, wire-speed anomaly detection by performing feature extraction and AI-based inference directly at the data plane, thereby identifying subtle attack patterns with reduced latency without compromising network throughput.
@@ 6.3 Optimized Traffic Management and Quality of Service
AI-optimized in-network computation is essential for dynamic and intelligent traffic management and Quality of Service (QoS) mechanisms [38]. AI models can be embedded into network devices to make necessary decisions about resource allocation, traffic routing, and prioritization. These decisions are based on application requirements and real-time network conditions [30]. In addition, AI tools can be used to optimize forwarding rule storage, enhance content caching, and anticipate user needs for better load balancing. To increase throughput and guarantee the required bandwidth, in-network ML can help with dynamic network path selection and service type prioritization. The ability to perform early-stage traffic classification in the data plane allows for the embedding of Service Level Objectives (SLOs) into packet headers, enabling stateless QoS provisioning [38].